Status Report

NOAA GOES SATOPS Morning Report 28 June 2001

By SpaceRef Editor
June 28, 2001
Filed under , ,


No change in the status of the POES

NOAA-15 rev 16229 / F at 1645Z: Multiple HIRS Flags. The NOAA-15
HIRS filter wheel
motor appears to be acting up again. There was a current spike up
to ~300 mA at ~1310z and
associated period monitor activity and loss of synchronization.
Things settled down (although
not to pre-1300z levels) until ~1435z
when another, wider, current spike (again ~300 mA) occurred,
again with loss of sync and
period monitor response. As a result of the erratic behavior, the
HIRS filter motor heater was
turned on yesterday at ~1825z (rev 16230). The temperature and
current rose for ~45
minutes, when the temperature leveled out, and the current
dropped somewhat, but was still
noisy and erratic. The current rose slightly again at ~2315z, but
both the current and
temperature began to fall (~linearly) at ~0100z. At ~0330z, the
current began jumping
between 2 levels (~260 mA – ~280 mA), a signature seen last
summer. Also at 0330z, the
period monitor and sync began to settle out. The latest data
shows the current reasonably
steady at ~260mA, but noisier than prior to this latest anomaly.
The period monitor is varying
~+/- 3 counts, and the filter wheel seems to be in sync. The plan
is to monitor the telemetry for
the rest of the day and to turn the filter motor heater off
tomorrow morning if things continue to
improve (or remain in their present state). The telemetry will be
monitored tomorrow and
through the weekend, with the heater ‘on’ kept as a contingency

POES operations were nominal over the past 24 hours.


No change in the status of the DMSP spacecraft.

OPS49 rev 7958 at Pogo had a 3367 spacecraft second / 84.175
playback second Cat I data
loss declared due to misconfiguration of ground system at SOCC.

DMSP operations were nominal over the past 24 hours.


No change in the status of the GOES spacecraft.

At 1941z and 1958z on June 27, GOES-8 frame breaks were observed
and attributed to RFI (unknown source).

At 0038z on June 28, a GOES-10 frame break was observed and
attributed to high BER

GOES operations were nominal over the past 24 hours.

Questions or comments to

SpaceRef staff editor.